# How connection probability shapes fluctuations of neural population dynamics

### v1

@@ Abstract
Mean-field models of neuronal populations in the brain have proven extremely useful to understand network dynamics and response to stimuli, but these models generally lack a faithful description of the fluctuations in the biologically relevant case of finite network size and connection probabilities $p<1$ (non-full connectivity). To gain insight into the different fluctuation mechanisms underlying the neural variability of populations of spiking neurons, we derive and analyze a stochastic mean-field model for finite-size networks of Poisson neurons with random, non-full connectivity, external noise and disordered mean inputs. We treat the quenched disorder of the connectivity by an annealed approximation that enables a reduction to a low-dimensional closed system of coupled Langevin equations for the mean and variance of the neuronal membrane potentials as well as a variable capturing finite-size fluctuations arising specifically in the case $p<1$. Comparing to microscopic simulations, we find that the mesoscopic model describes the fluctuations and nonlinearities well and outperforms previous mesoscopic models that neglected the recurrent noise effect caused by the non-full connectivity. This effect can be analytically understood by a softening of the effective nonlinearity and the multiplicative character of finite-size spiking noise. The mesoscopic theory shows that quenched disorder can stabilize the asynchronous state, and it correctly predicts large quantitiative and non-trivial qualitative effects of connection probability on the variance of the population firing rate and its dependence on stimulus strength. Our theory thus elucidates how disordered connectivity shapes nonlinear dynamics and fluctuations of neural populations at the mesoscopic scale and showcases a useful mean-field method to treat non-full connectivity in finite-size, spiking neural networks.
@@ I Introduction
Biological systems often exhibit significant fluctuations and variability in their dynamics, in contexts ranging from gene expression to neural activity in the brain. The possible roles of fluctuations for the dynamics and functions of such systems have thus received increasing attention [1, 2, 3, 4, 5, 6, 7]. An important tool to understand the collective dynamics of biological systems consisting of many interacting units are low-dimensional descriptions of the effective dynamics at the population level. Taking fluctuations properly into account can be necessary to establish a correct understanding of the system and can provide crucial insight into the origins of observed biological variability. At the population level, fluctuations may then be accounted for by stochastic low-dimensional models. In the context of living systems, examples range from Langevin equations for gene regulation [8] to coupled molecular motors [9, 10, 11] to ecology [12].
@@
In neuroscience, neural-mass models (also called firing-rate or neural-population models) such as the Wilson-Cowan model [13] have been highly successful in describing various phenomena in cortical circuits such as oscillations [14], multistability [15], and nonlinear reponse properties [16]. Stochastic versions of neural-mass models have been critical for understanding fluctuating neural population dynamics such as metastability [17, 18, 15], stochastic oscillations [19], and response properties of cortical variability [20, 21]. While neural-mass models and their stochastic variants stand out due to their mathematical tractability, they are often phenomenological, lacking a clear link to an underlying network of spiking neurons at the microscopic level.
Neural population models that maintain a link to biophysical properties can be obtained through mean-field modeling [22, 23, 24, 25, 26, 27, 28], where the macroscopic or mesoscopic dynamics is analytically derived from a microscopic model. The resulting mean-field models are amenable to mathematical analysis [29, 30, 31, 32] or may enable efficient and accurate simulations of large spiking neural networks [33, 34, 35, 36]. In current mean-field models, however, the treatment of fluctuations is typically based on the (implicit) assumptions of either infinitely large network size or homogeneous, full (all-to-all) connectivity. These assumptions are often not applicable to biologically realistic recurrent neural networks. For example, in rat barrel cortex, the number of neurons per cell type in a given layer of a column has been counted to be on the order of $N\sim 100$ to $N\sim 1000$ neurons per cell type [37] – a population size for which finite-size fluctuations cannot be neglected. Furthermore, the local connection probabilities $p$ in mouse visual cortex vary on a wide range ($0\lesssim p\lesssim 0.7$) with low average ($p\sim 0.1$) and larger connection probabilities ($p\sim 0.5$) for similarly tuned neurons [38].
@@
Both finite network sizes and connection probabilities $<1$ cause intrinsically generated (endogenous) fluctuations of the recurrent, synaptic input currents, and may thus strongly affect the nonlinear population dynamics. However, their distinct and combined effects are still poorly understood. Moreover, these intrinsic fluctuations act as multiplicative noise whose strength increases with population firing rates. On the other hand, neural variability in the cortex often decreases upon stimulation [39] – a phenomenon that could be explained, among others, with a model, where variability is generated by an external, additive noise [21]. To clarify the role of intrinsic fluctuations for such cortical phenomena under biologically realistic neuron numbers and connection probabilities requires a mean-field model that accounts for these different origins of fluctuations.
A mesoscopic theory for finite-size spiking neural networks providing accurate population models with fluctuations has recently been developed for generalized integrate-and-fire neurons with escape noise [35, 32]. This theory suggests a multiscale modeling framework for efficient forward [35] and inverse [40, 41] modeling of cortical circuits based on realistic neuron dynamics [42, 43]. However, this and other mean-field theories [44, 45, 35, 46, 47, 48, 32] for finite-size neural networks are based on homogeneous, fully connected networks ($p=1$, “global coupling”). Even though the full connectivity may be regarded as an effective, mean-connectivity approximation of a non-fully connected network [49, 35, 50], the validity of such approximation remains unclear. In particular, in a homogeneous fully connected network, the fluctuations caused by a finite network size $N<\infty$ are common to all neurons because neurons receive identical recurrent inputs when neurons are connected all-to-all. On the population level, these “coherent” fluctuations lead to stochastic population equations with finite-size noise of order $1/\sqrt{N}$. Clearly, in random networks with connection probability $p<1$, different neurons no longer receive identical synaptic inputs, and therefore the implicit assumption of perfectly coherent noise made in the effective, mean-connectivity approximation cannot hold fully true.
@@
A mean-field description of non-fully connected, random networks ($p<1$) is difficult because it requires the treatment of fluctuations resulting from a “quenched” (i.e. temporally constant) random synaptic connectivity. A classical approach to deal with random networks is to consider the macroscopic limit $N\to\infty$. Depending on the concrete limiting procedure and network architecture, the recurrent fluctuations seen by different neurons either vanish [22, 30, 31, 28] or become perfectly independent (“incoherent”) in this limit [51, 52, 53, 54, 55] unless correlations are induced externally. A famous example is the mean-field theory for sparsely connected networks by Brunel and co-workers [29, 51]. In this theory, the neural population dynamics is deterministic and given by a nonlinear Fokker-Planck equation. The effect of incoherent fluctuations of the recurrent synaptic input appears as an additional contribution to the diffusion coefficient in this equation. Although this approach neglects temporal correlations of the spike trains caused by the quenched disorder, the nonlinear Fokker-Planck equation has been successful to capture the influence of intrinsic fluctuations on the nonlinear population dynamics (for a treatment of temporal correlations, see e.g. [56]). Using this theory, intrinsic fluctuations have been shown to crucially shape nonlinear response properties [57], network oscillations [29, 58] and multistability [59].
@@
Again, even though mean-field theories for $N\to\infty$ have been used as simple mean-field descriptions for real neural networks of finite size (see e.g. [60]), they cannot describe finite-size fluctuations at the population level. Furthermore, finite-size networks are necessarily outside the sparse limit (i.e. $p>0$) which entails a non-vanishing probability of shared recurrent inputs among neurons. Therefore the assumption of perfectly incoherent noise in the mean-field theory cannot hold fully true for $N<\infty$. While finite-size extensions of the mean-field theory for sparsely connected networks have been proposed in [29, 51, 61], low-dimensional neural population models that reveal and take into account the distinct effects of coherent and incoherent fluctuations are currently lacking.
@@
In this paper, we derive and analyze a simple bottom-up mean-field model that describes incoherent and coherent recurrent fluctuations for arbitrary connection probabilities $p\in[0,1]$ in a finite-size network. To this end, we study a network of Poisson neurons with “annealed” random connectivity [62, 29] as an approximation to a corresponding network with “quenched” random connectivity. The annealed approximation neglects temporal correlations of incoherent fluctuations but captures the effect of non-full connectivity and finite network size surprisingly well. Importantly, the annealed approximation enables a simple mean-field description of the fluctuations in form of a three-dimensional Langevin equation. We compare this model with a naive mean-field model in which the incoherent fluctuations due to random connectivity are neglected and which precisely corresponds to the ad hoc mean-field approximation used in previous models [49, 35, 50]. While the focus of the paper is the Langevin dynamics for a homogeneous population with quenched or annealed random connectivity and common external noise, the theoretical framework can be extended to heterogeneous populations, where the resting potentials or external currents of the neurons exhibit quenched Gaussian disorder. In summary, we present a simple theoretical framework for the effects of various sources of noise and quenched disorder on the neural population dynamics at the mesoscopic scale ($1\ll N<\infty$).
@@ II.1 Quenched network model
At the microscopic level, we consider a random network of $N$ interacting Poisson neurons (also called nonlinear Hawkes processes). In this model, neurons fire spikes stochastically with conditional intensity $\lambda_{i}(t)=\phi(h_{i}(t^{-}))$, $i=1,\ldots,N$, where $h_{i}(t)$ is the input potential of neuron $i$ and $\phi$ is a non-negative hazard function. The dynamics of the input potentials is given by a system of coupled first-order equations with delay,
@@
$i=1,\dotsc,N$. Here, $\tau$ is the time constant of the low-pass filter dynamics, $\mu(t)$ is an external drive and $\dot{Z}_{i}(t)=\sum_{l}\delta(t-t_{i,l})$ is the spike train of neuron $i$ with $\{t_{i,l}\}$ being the individual spike times. The presence or absence of synaptic connections between neurons are described by the random adjacency matrix $[a_{ij}]$ with fixed in-degree $C=\sum_{j=1}^{N}a_{ij}$ and elements $a_{ij}\in\{0,1\}$ that are marginally Bernoulli distributed with mean $\left\langle a_{ij}\right\rangle=p=C/N$ representing the connection probability. This random connectivity can be constructed by choosing for each neuron $C$ presynaptic neurons randomly, and the resulting connectivity is fixed (“quenched”) in time. Furthermore, $d$ is the synaptic delay and $w=CJ$ is the total coupling strength with $J$ being the efficacy of a single synapse (i.e. $J/\tau$ is the jump size in millivolt of the postsynaptic potential caused by a single presynaptic spike). Since we focus on inhibitory networks in this paper, we assume that $J<0$.
@@ II.2 Mean-connectivity network
Figure 1: Response of the population rate variance to a change of the mean stimulus strength in the original, quenched random network (a) and in a corresponding fully-connected, mean-connectivity network that only accounts for the mean synaptic weight (b). Top: Noisy external stimulus $\mu(t)$ whose mean increases at $t=0.08$ s in a step-like fashion from $\mu_{0}=28$ mV to $\mu_{0}=50$ mV; the noise strength $\sigma_{\text{ext}}^{2}=1$ mV2 is kept fixed. (a) Schematic of a quenched network with fixed adjacency matrix $a_{ij}$ (left) and the trial-averaged variance of the population firing rate as a function of time (right) for different network sizes (green line: $N=5\cdot 10^{4}$, blue line: $N=10^{3}$, in both cases $C=100$). (b) Same for a mean-connectivity network, Eq. (5), as an ad hoc approximation of the quenched network, where the quenched connectivity $a_{ij}$ is replaced by the average connectivity $\langle a_{ij}\rangle$. Note the different scaling of the y-axis in a) and b). Thus, while the mean stationary population rates after the step (as indicated in the right panels) are equal in both models, the variances of the population rate strongly differ.
[…]
@@
The replacement of all elements of the adjacency matrix $a_{ij}$ with their mean $p$ has been used ad hoc also for finite network size $N<\infty$ [49, 35, 50]. However, this ad hoc approach is no longer exact for finite $N$ but must be regarded as an approximation that retains the mean synaptic connectivity $\left\langle J_{ij}\right\rangle$ but ignores all higher-order cumulants of the random variables $J_{ij}$. In the following, we will refer to this effective, fully connected network as the mean-connectivity network, whereas the original model, Eq. (1), will be referred to as the quenched network. Replacing the connection strengths $w_{ij}$ with their means $w$, we obtain from Eq. (1) the dynamics of the mean-connectivity network:
@@
where $dZ_{j}(t)\sim\text{Pois}(\phi(h_{j}(t))dt)$ are the conditionally independent Poisson increments of the spike count of neuron $j$. The mean-connectivity network of Poisson neurons, Eq. (5), is a microscopic model that admits an exact reduction to a mesoscopic model for the total spike count $dZ(t)=\sum_{i=1}^{N}dZ_{i}(t)$. To see this, we note that the sum of independent Poisson numbers in Eq. (5) is a Poisson number with mean $Nr(t-d)dt$, where
@@
where now $dZ(t)\sim\text{Pois}(N\phi(\bar{h}(t))dt)$ describes the increment of the total spike count of the population. Equation (7) describes the mesoscopic dynamics as it only involves the population activity $dZ(t)/(Ndt)$.
Simulations of the quenched and the mean-connectivity network yield similar results for the stationary mean firing rates and mean input potentials. However, the mean-connectivity network may differ drastically compared to the quenched network when it comes to the non-stationary response (linear response function) and the fluctuations (second-order statistics). To illustrate the failure of the mean-connectivity approximation, let us consider the variance of the stochastic population rate $r(t)$. This variance represents the part of the mesoscopic neural variability (variance of the population activity $A_{N}(t,\Delta t)$) that is caused by the rate variability rather than the Poisson spiking noise. It can be measured from population activity data (see Appendix, Sec. B), or directly computed from a simulated time series of $r(t)$ in our model. A non-vanishing variance of the population rate can occur for two reasons in our model: first, because of intrinsic finite-size fluctuations, and second, because of externally injected common noise. Here, we model the common external noise as a Gaussian white noise (more precisely, $\mu(t)=\bar{\mu}(t)+\sqrt{\tau\sigma_{\text{ext}}^{2}}\hat{\zeta}(t)$, where $\bar{\mu}(t)$ is the mean stimulus and $\hat{\zeta}(t)$ is a standard Gaussian white noise process).
Interestingly, simulations of the mean-connectivity network (Eq. (5) or (7)) strongly overestimate the population-rate variance of the quenched network by more than one order of magnitude (Fig. 1, note the different scale of the y axes). In addition to this large quantitative difference, we also observe a qualitatively different response to stimulus changes. When the mean stimulus strength is increased from $28$ mV to $50$ mV, the variance of the population rate exhibits a small but significant decrease in the quenched network (Fig. 1a, right). This suppression of variability is in marked contrast to the prediction by the mean-connectivity network, which exhibits instead an increase in the variance (Fig. 1b, right). We stress that the marked differences concern the second-order statistics but not the first-order statistics: The stationary mean population rates $\left\langle r\right\rangle$ are the same in both networks as indicated in Fig. 1a,b.

### v2

# How random connectivity shapes the fluctuating dynamics of finite-size neural populations

@@ Abstract
Mesoscopic models of finite-size neuronal populations are crucial to understand the dynamics of neural networks in the brain, especially their fluctuations and response to stimuli. However, current theories to derive such models are based on the assumption of homogeneous all-to-all (full) connectivity, neglecting the variance in the connectivity of biologically realistic networks with connection probabilities $p<1$ (non-full connectivity). To gain insight into the different fluctuation mechanisms underlying the neural variability of populations of spiking neurons, we derive and analyze a stochastic mean-field model for finite-size networks of Poisson neurons with random connectivity (including non-full connectivity), external noise and disordered mean inputs. We treat the quenched disorder of the connectivity by an annealed approximation enabling a doubly stochastic description of synaptic inputs for finite network size. A further reduction leads to a low-dimensional closed system of coupled Langevin equations for the mean and variance of the neuronal membrane potentials as well as a variable capturing finite-size fluctuations arising specifically in the case of connectivity disorder. Comparing to microscopic simulations, we find that the mesoscopic model describes the fluctuations and nonlinearities well and outperforms previous mesoscopic models that neglected the variance in the connectivity. The joint effect of connectivity disorder and finite network size can be analytically understood by a softening of the effective nonlinearity and the multiplicative character of spiking noise. The mesoscopic theory shows that quenched disorder can stabilize the asynchronous state, and it correctly predicts large quantitative and non-trivial qualitative effects of connection probability on the variance of the population firing rate and its dependence on stimulus strength. Our theory thus elucidates how disordered connectivity shapes nonlinear dynamics and fluctuations of neural populations at the mesoscopic scale and showcases a useful mean-field method to treat random connectivity in finite-size, spiking neural networks.
@@ I Introduction
Biological systems often exhibit significant fluctuations and variability in their dynamics, in contexts ranging from gene expression to neural activity in the brain. The possible roles of fluctuations for the dynamics and functions of such systems have thus received increasing attention [1, 2, 3, 4, 5, 6, 7]. An important tool to understand the collective dynamics of biological systems consisting of many interacting units are low-dimensional descriptions of the effective dynamics at the mesoscopic or macroscopic scale. Taking fluctuations properly into account can be necessary to establish a correct understanding of the system and can provide crucial insight into the origins of observed biological variability. At the mesoscopic scale, fluctuations naturally arise from the finite number of underlying units. These “finite-size” fluctuations may then be accounted for by stochastic low-dimensional models. In the context of living systems, examples range from Langevin equations for gene regulation [8] to coupled molecular motors [9, 10, 11] to ecology [12].
@@
In neuroscience, neural-population models (also known as firing-rate or neural-mass models) such as the Wilson-Cowan model [13] have been highly successful in describing various cortical phenomena such as oscillations [14], multistability [15], and nonlinear response properties [16]. Stochastic versions of neural-population models have been critical for understanding fluctuating neural population dynamics such as metastability [17, 18, 15], stochastic oscillations [19], and response properties of cortical variability [20, 21]. Despite their mathematical tractability, however, these models are often phenomenological, lacking a clear link to an underlying finite-size network of spiking neurons.
A link to biophysical properties is maintained through bottom-up mean-field approaches [22, 23, 24, 25, 26, 27, 28], where the macroscopic or mesoscopic dynamics is analytically derived from a microscopic model. The resulting mean-field dynamics are amenable to mathematical analysis [29, 30, 31, 32, 33] and enable efficient and accurate simulations of large spiking neural networks [34, 35, 36, 37]. Recently, a bottom-up theory for the mesoscopic dynamics of finite-size populations of generalized integrate-and-fire neurons with escape noise has been developed [36, 33]. It suggests an efficient multiscale modeling framework of cortical circuits [36, 38, 39] grounded in realistic neuron dynamics [40, 41]. However, this and other mean-field theories [42, 43, 36, 44, 45, 46, 33] for finite-size neural networks typically assume homogeneous, all-to-all (full) connectivity. This assumption can be regarded as an approximation of a non-fully connected random network by its homogeneous mean connectivity while neglecting its quenched variability [47, 36]. On the other hand, the role of random connectivity for intrinsically generated fluctuations of synaptic inputs has been extensively studied using dynamical mean-field theories [48, 49, 29, 30, 50, 51, 52, 53], albeit only in the limit of infinitely large networks.
Both assumptions–homogeneous full connectivity ignoring quenched variability and infinite network size ignoring fluctuations–are severe limitations for describing biologically realistic, recurrent neural networks. For example, rat barrel cortex contains only on the order of $N\sim 100$ to $N\sim 1000$ neurons per cell type per layer [54] – a population size for which finite-size fluctuations can be significant, as shown in computational models [55, 56, 57, 36, 58, 59]. Furthermore, the local connection probabilities $p$ in mouse visual cortex vary on a wide range ($0\lesssim p\lesssim 0.7$) with low average ($p\sim 0.1$) and larger connection probabilities ($p\sim 0.5$) for similarly tuned neurons [60]. How such pronounced deviations from fully connected networks ($p=1$) affect the mesoscopic dynamics of finite-size neural populations is still poorly understood.
@@
Both finite population size ($N<\infty$) and quenched random connectivity (e.g. $p<1$) cause intrinsically generated (endogenous) fluctuations of the recurrent, synaptic input currents. These intrinsic fluctuations act as multiplicative noise whose strength increases with population firing rates. On the other hand, neural variability in the cortex often decreases upon stimulation [61] – a phenomenon that could be explained, among others, with a model where variability is generated by an external, additive noise [21]. To clarify the role of intrinsic fluctuations for such cortical phenomena under biologically realistic neuron numbers and connection probabilities requires a mean-field theory that accounts for these different origins of fluctuations.
In a homogeneous, fully connected network, as arising under a mean-connectivity approximation, the input fluctuations are common to all neurons. This is because all neurons receive identical recurrent inputs and thus experience the same finite-size fluctuations of the population activity. These “coherent” fluctuations lead to stochastic population equations with finite-size noise of order $1/\sqrt{N}$. In contrast, in random networks with connection probability $p<1$, neurons no longer receive identical synaptic inputs, and hence input fluctuations become partially decorrelated. Consequently, the assumption of perfectly coherent fluctuations no longer holds, raising questions about the validity of the mean-connectivity approximation for finite-size networks.
@@
A mean-field description of random networks (including non-fully connected networks, $p<1$) requires the treatment of fluctuations resulting from a “quenched” (i.e. temporally constant) random synaptic connectivity. Previous approaches to random networks considered the macroscopic limit $N\to\infty$. Depending on the concrete limiting procedure and network architecture, the recurrent fluctuations seen by different neurons either vanish [22, 31, 32, 28] or become perfectly independent (“incoherent”) in this limit [48, 50, 29, 30, 51, 52, 53] unless correlations are induced externally. A famous example is the mean-field theory for sparsely connected networks by Brunel and co-workers [29, 30]. In this theory, the neural population dynamics is deterministic and given by a nonlinear Fokker-Planck equation. The effect of incoherent fluctuations of the recurrent synaptic input caused by the random connectivity appears as an additional contribution to the diffusion coefficient in this equation. Although this approach relies on an annealing approximation that neglects temporal correlations in the spike trains caused by the quenched disorder, the nonlinear Fokker-Planck equation has been successful to capture the influence of intrinsic fluctuations on the nonlinear population dynamics (for a treatment of temporal correlations, see e.g. [62]). Using this theory, intrinsic fluctuations have been shown to crucially shape nonlinear response properties [63], network oscillations [29, 64] and multistability [56].
@@
Even though dynamical mean-field theories for $N\to\infty$ have been used for real neural networks of finite size (see e.g. [65]), they cannot describe finite-size fluctuations at the population level. Furthermore, finite-size networks are necessarily outside the sparse limit (i.e. $p>0$) which entails a non-vanishing probability of shared recurrent inputs among neurons. Therefore the assumption of perfectly incoherent noise in dynamical mean-field theories cannot hold fully true for $N<\infty$. While finite-size extensions of the mean-field theory for sparsely connected networks have been proposed in [29, 30, 66], low-dimensional neural population models that reveal and take into account the distinct effects of coherent and incoherent fluctuations are currently lacking.
@@
In this paper, we consider finite-size networks with random connectivity, allowing arbitrary connection probabilities $p\in[0,1]$. For these networks, we derive and analyze a simple mesoscopic dynamics that describes incoherent and coherent recurrent fluctuations. To this end, we study a network of Poisson neurons using an annealed approximation [67, 29] of the quenched random connectivity. The annealed approximation neglects temporal correlations of incoherent fluctuations but captures the effects of random connectivity and finite network size surprisingly well. In contrast to the case $N\to\infty$ of previous theories [29, 30], we need to treat the fluctuating input as a doubly stochastic process to account for both finite $N$ and random connectivity. This description eventually yields a three-dimensional Langevin equation for the mesoscopic dynamics. We compare this dynamics with a naive mean-field theory in which the variance of the random connectivity is neglected, corresponding to the ad hoc mean-field approximation used in previous studies [47, 36]. While the focus of the paper is the Langevin dynamics for a homogeneous population with quenched or annealed random connectivity and common external noise, the theoretical framework can be extended to heterogeneous populations, where the resting potentials or external currents of the neurons exhibit quenched Gaussian disorder. In summary, we present a simple theoretical framework for the effects of various sources of noise and quenched disorder on the neural population dynamics at the mesoscopic scale ($1\ll N<\infty$).
@@ II.1 Quenched network model
At the microscopic level, we consider a random network of $N$ interacting Poisson neurons (also called nonlinear Hawkes processes). While our theory will be developed for general random connectivity with finite variance, we focus on the specific example of binary, non-full connectivity to study the effect of connection probabilities $p<1$. The Poisson neurons fire spikes stochastically with conditional intensity $\lambda_{i}(t)=\phi(h_{i}(t^{-}))$, $i=1,\ldots,N$, where $h_{i}(t)$ is the input potential of neuron ii and $\phi$ is a non-negative hazard function. The dynamics of the input potentials is given by a system of coupled first-order equations with delay,
@@
$i=1,\dotsc,N$. Here, $\tau$ is the time constant of the low-pass filter dynamics, $\mu(t)$ is an external drive and $\dot{Z}_{i}(t)=\sum_{l}\delta(t-t_{i,l})$ is the spike train of neuron $i$ with $\{t_{i,l}\}$ being the individual spike times. The presence or absence of synaptic connections between neurons are described by the random adjacency matrix $a_{ij}]$ with fixed in-degree $C=\sum_{j=1}^{N}a_{ij}$ and elements $a_{ij}\in\{0,1\}$ that are marginally Bernoulli distributed with mean $\left\langle a_{ij}\right\rangle=p=C/N$ representing the connection probability. This random connectivity can be constructed by choosing for each neuron $C$ presynaptic neurons randomly, and the resulting connectivity is fixed (“quenched”) in time. Furthermore, $d$ is the synaptic delay and $w=CJ$ is the total coupling strength with $J$ being the efficacy of a single synapse (i.e. $J/\tau$ is the jump size in millivolt of the postsynaptic potential caused by a single presynaptic spike). Since we focus on inhibitory networks in this paper, we assume that $J<0$.
@@ II.2 Mean-connectivity network
Figure 1: Response of the population rate variance to a change of the mean stimulus strength in the original, quenched random network (a) and in a corresponding fully-connected, mean-connectivity network that only accounts for the mean synaptic weight (b). Top: Noisy external stimulus $\mu(t)$ whose mean increases at $t=0.08$ s in a step-like fashion from $\mu_{0}=28$ mV to $\mu_{0}=50$ mV; the noise strength $\sigma_{\text{ext}}^{2}=1$ mV2 is kept fixed. (a) Schematic of a quenched network with fixed adjacency matrix $a_{ij}$ (left) and the trial-averaged variance of the population firing rate as a function of time (right) for different network sizes (green line: $N=5\cdot 10^{4}$, blue line: $N=10^{3}$, in both cases $C=100$). (b) Same for a mean-connectivity network, Eq. (5), as an ad hoc approximation of the quenched network, where the quenched connectivity $a_{ij}$ is replaced by the average connectivity $\langle a_{ij}\rangle$. While the mean stationary population rates after the step are equal in both models (measured rates are provided in the right panels), the variances of the population rate strongly differ (note the different scaling of the $y$-axis). Other parameters: $\beta=5$ mV-1, w=1w=-1 mVs.
A major obstacle in deriving population models analytically is the quenched randomness of the connectivity $J_{ij}:=Ja_{ij}$. A standard mean-field approach to tackle this problem is to neglect fluctuations in the recurrent synaptic input by scaling the synaptic weights inversely proportional to the network size, $J_{ij}=w_{ij}/N$, where the coupling strength $w_{ij}$ is assumed to be of order $1$. Then, in the limit $N\to\infty$, the synaptic input $N^{-1}\sum_{j}w_{ij}\dot{Z}_{j}$ converges to the deterministic input $\left\langle w_{ij}\right\rangle\langle\dot{Z}_{j}\rangle$ [27, 28], where $\left\langle\cdot\right\rangle$ denotes the average over both quenched disorder and Poisson noise. In our case, this corresponds to a network where the quenched connectivity $w_{ij}$ (Fig. 1a, left) is replaced by the mean connectivity $\left\langle w_{ij}\right\rangle=NJ\left\langle a_{ij}\right\rangle{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=CJ}=w$ (Fig. 1b, left). Importantly, the mean connectivity effectively yields a fully connected network with homogeneous coupling $w$. In the limit $N\to\infty$, the synaptic input is thus not affected by the variance
@@
The replacement of all elements of the adjacency matrix $a_{ij}$ with their mean $p$ has been used ad hoc also for finite network size $N<\infty$ [47, 36]. However, this ad hoc approach is no longer exact for finite $N$ but must be regarded as an approximation that retains the mean synaptic connectivity $\left\langle J_{ij}\right\rangle{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=pJ}$ but ignores all higher-order cumulants of the random variables $J_{ij}$. In the following, we will refer to this effective, fully connected network as the mean-connectivity network, whereas the original model, Eq. (1), will be referred to as the quenched network (Table 1). Replacing the connection strengths $w_{ij}$ with their means $w$, we obtain from Eq. (1) the dynamics of the mean-connectivity network:
@@
where $dZ_{j}(t)\sim\text{Pois}(\phi(h_{j}(t))dt)$ are the conditionally independent Poisson increments of the spike count of neuron $j$. The mean-connectivity network of Poisson neurons, Eq. (5), is a microscopic model that admits an exact reduction to a mesoscopic model for the total spike count $dZ(t)=\sum_{i=1}^{N}dZ_{i}(t)$. To see this, we note that the sum of independent Poisson numbers in Eq. (5) is a Poisson number with mean Nr(td)dtNr(t-d)dt, where
@@
where now $dZ(t)\sim\text{Pois}(N\phi(\bar{h}(t))dt)$ describes the increment of the total spike count of the population. Equation (7) describes the mesoscopic dynamics as it only involves the population activity $A_{N}(t,\Delta t)=\int_{t}^{t+\Delta t}dZ(t)/(N\Delta t)$.
Simulations of the quenched and the mean-connectivity network yield similar results for the stationary mean firing rates and mean input potentials. However, the mean-connectivity network may differ drastically compared to the quenched network when it comes to the non-stationary response (linear response function) and the fluctuations (second-order statistics). To illustrate the failure of the mean-connectivity approximation, let us consider the variance of the stochastic population rate $r(t)$. This variance represents the part of the mesoscopic neural variability (variance of the population activity $A_{N}(t,\Delta t)$) that is caused by the rate variability rather than the Poisson spiking noise. It can be measured from population activity data (see Appendix, Sec. C), or directly computed from a simulated time series of $r(t)$ in our model. A non-vanishing variance of the population rate can occur for two reasons in our model: first, because of intrinsic finite-size fluctuations, and second, because of externally injected common noise. Here, we model the common external noise as a Gaussian white noise (more precisely, $\mu(t)=\bar{\mu}(t)+\sqrt{\tau\sigma_{\text{ext}}^{2}}\hat{\zeta}(t)$, where $\bar{\mu}(t)$ is the mean stimulus and $\hat{\zeta}(t)$ is a standard Gaussian white noise process).
Interestingly, simulations of the mean-connectivity network (Eq. (5) or (7)) strongly overestimate the population-rate variance of the quenched network by more than one order of magnitude (Fig. 1, note the different scale of the y axes). We stress that this difference is not an effect of a difference in firing rates: The stationary mean population rates $\left\langle r\right\rangle$ are the same in both networks (e.g., the numerical values of the firing rates at high stimulus amplitudes are approximately equal as indicated in Fig. 1a,b). In addition to this large quantitative difference, we also observe a qualitatively different response to fast stimulus changes when the network is large ($N=5\cdot 10^{4}$, green lines in Fig. 1): When the mean stimulus strength jumps from $28$ mV to $50$ mV, the variance of the population rate exhibits a small but significant decrease in the quenched network (Fig. 1a, right). This suppression of variability is in marked contrast to the prediction by the mean-connectivity network, which exhibits instead an increase in the variance (Fig. 1b, right). This qualitatively different behavior of the population-rate variance in response to a step stimulus is, however, not observed for smaller network size ($N=10^{3}$, but otherwise same parameters and similar firing rates; blue lines in Fig. 1).